OpenStudy (anonymous):

implicit differentiation y = sin(7xy) show steps please

OpenStudy (anonymous):

\[y'=7\cos(7xy)\times(y+xy')\] is a start

OpenStudy (anonymous):

only not obvious part is that the derivative of \[xy\] requires the product rule, since you are thinking of y as a function of x. like taking the derivative of \[xf(x)\] using the product rule. you would get \[(xf(x))'=f(x)+xf'(x)\] but here you write \[y+xy'\]

OpenStudy (anonymous):

then algebra to solve for \[y'\]

OpenStudy (anonymous):

whats is wrong with this site, i am trying to post and answer and it freezes on me and then i have to refresh and i lose everthing i wrote

OpenStudy (anonymous):

very very slow tonight

OpenStudy (anonymous):

yeah its been slow for a couple of nights now

OpenStudy (anonymous):